OpenStudy (anonymous):

Integration via U-sub: Integral (sqrt(x)+1)^4/2sqrt(x)

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ (\sqrt{x}+1)^{4} }{ 2\sqrt{x} }\]

OpenStudy (anonymous):

I'm a bit unsure of how to do this one. So far I have this: u = x+1 du/sx =1 => du=1dx Would I have to solve for x in u to get: x=u-1 So would my integral with u sub look like this: \[\int\limits_{}^{}\frac{ u^4 }{ u-1 }\]

OpenStudy (bradely):

u=sqrtx+1 du/dx =1/2sqrtx du=dx/2sqrtx int(udu) =u^2/2+C= (sqrtx+1)^2/2 Source: http://www.mathskey.com/question2answer/